Grade 12Physical Chemistry

The wavelength of a neutron with a translator kinetic energy equal to KT at 300 kelvin is?

Nikunj Gour
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
you have E= mv^2 /2 

so v^2 = 2E/m m of neutron =1.67 10^(-27) kg 

so v^2 = 2E/m E = kT = 1.38 10^-23*300 =4.14 10^(-21) J 

v^2 = 2*4.14 10^(-21)/1.67 10^(-27) =4.96 10^6 m^2s^-2 

v=2.227 10^3 m/s 

And 
lambda = 6.6210^(-34)/(2.22710*3)*1.67 10^(-27) 
lambda=1.78 10^(-10)m =0.178 nm